OpenStudy (anonymous):

Assume g(x) = (x-1)/(x+2). Find a number such that g(b) = 5

OpenStudy (campbell_st):

well won't you have \[5 = \frac{(x -1)}{(x+2)}\] multiply both sides of the equation by (x +2) which gives 5(x +2) = x - 1 then just distribute and solve for x.

OpenStudy (anonymous):

Aha, thanks so much. Been a few years since I've taken a math course so a lot of things I used to be great at, including questions like this have leaked out of my head. Beyond frustrating.